The time required to drive from Duluth to Fargo varies inversely with the speed of the car. If the trip takes 6.25 hours at 40 mph, what speed is required to make the trip in 4 hours?
A) 6.4 mph B) 0.625 mph C) 25.6 mph D) 8.44 mph E) 62.5 mph

Answer by richwmiller(17219)

r*t=d
t=d/r
6.25*40=x*4
6.25*10=x
62.5=x
E) 62.5 mph